A valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Barják Tamás  +  +  +  +  −  +  I
Barkován Anett  +  +  +  −  +  −  I
Bihari Szabina  +  +  ±  +  +  ±  I
Bócsi Judit  +  ±  ±−  ±  +  −  I
Bodnár Zsuzsanna  +  +  ±  −  +  +  I
Breszkócs Péter Lajos  ±  −  −  ±  −  −  N
Flaskár Ibolya  +  −  −  −  +  ±  I
Gajdics László  +  +  −  −  +  +  I
Hegedűs Judit  ±  +  −  +  +  ±  I
Herczik Bettina  ±  +  −  +  ±  +  I
Husiová Anikó  −  −  −  −  −  −  N
Karakas Éva  ±  +  ±  −  ±  −  I
Kiss Annamária  +  +  −  −  −  −  N
Kovács György  −  −  −  −  −  −  N
Kovács Szilvia  ±  −  −  −  −  ±  N
Kovács Viktor  +  +  ±  −  −  −  I
Kovács Zoltán  −  −  −  −  −  −  N
Kovácsné, Sz. Renáta  ±  ±  −  −  −  −  N
Kozma Nikoletta  +  +  −  −  +  −  I
Kövér Ildikó  +  ±  −  −  +  −  I
Kurtos Adrienn  ±  +  −  +  +  −  I
Miklovszki Tamás  +  ±  +  −  +  −  I
Molnár Péter Miklós  +  ±  −  −  +  −  I
Németh Ildikó  +  +  ±  −  −  ±  I
Palcsesz Viktória  +  +  −  −  −  −  N
Péter Antal  +  +  −  −  −  −  N
Sárközi Borbála  +  +  −  −  ±  +  I
Sebestény Csilla  +  +  −  ±  −  −  I
Tóth Andrea  +  +  −  +  +  −  I
Tóth Eszter  +  +  +  +  −  +  I
Tóth István Márk  ±  −  −  −  ±  −  N
Vígh Csilla  +  +  +  +  −  +  I